# Non-Unitary Quantum Many-Body Dynamics using the Faber Polynomial Method

### v2

@@ 7 Conclusions
This study offers additional insights into the entanglement transition in quantum spin chains exhibiting Non-Hermitian or Liouvillian skin effects. In the Hatano-Nelson model with quantum jumps, we found that the area law behaviour of entanglement entropy remains for any nonzero $\gamma$, similar to the no-click limit. However, the sources are different: in the no-click limit, the area law stems from the single-particle skin effect, while in the monitored stochastic trajectory, it is due to the quantum jumps that relax the system into the fully polarised down spin-state, a product state. On the other hand, the dynamics with two-body jump operators, $L_{\ell}\propto S^{+}_{\ell}S^{-}_{\ell+1}$, allow for a non-equilibrium steady state with a magnetisation profile resembling the no-click limit of the Hatano-Nelson model. Furthermore, the average entanglement entropy still follows an area law for finite values of the ratio $\gamma/J$ in conditional dynamics. The measurement apparatuses effectively disentangle the state, suppressing the volume law that would otherwise be generated in unitary dynamics. This work extends previous studies that focused on the one-particle sector [105, 34].
Overall, our results support the utility and applicability of Faber polynomials in various research domains. This encompasses investigations into measurement-induced phase transitions, open quantum systems, and the exploration of purely non-unitary dynamics governed by non-Hermitian Hamiltonians. Faber polynomials can also be used for computing general spectral properties of non-Hermitian Hamiltonians, potentially replicating the role of Chebyshev polynomials within the Kernel polynomial method [64]. This would complement the existing Non-Hermitian Kernel polynomial method [113, 114], which still uses Chebyshev polynomials, but relies on hermitianisation techniques, at the cost of working in a vector space with twice the original dimension. Furthermore, the Faber polynomial approach could potentially be combined with MPS, similar to the developments already made in the Hermitian case  [115, 116, 117, 118].
@@ Appendix A Further Details on Faber Polynomials
$F_{n+1}(z)=zF_{n}(z)-\sum_{j=0}^{n}\gamma_{j}F_{n-j}(z)-n\gamma_{n},\quad n>0$ (49)
where $F_{0}(z)=1$. For our purposes, we are interested in using the Faber polynomials to approximate a given function of our non-Hermitian Hamiltonian, $f\left(\mathcal{H}\right)$. The domain $\mathcal{D}$, is defined by the spectrum of the Hamiltonian. Using Eq. 47, the expression in a Faber series is given by
$f(z):=\sum_{k=-\infty}^{+\infty}c_{k}F_{k}(z).$ (50)
@@
As stated in the main text, we perform this integral with $\rho=1$ by properly rescaling the Hamiltonian. Furthermore, we compute the Faber coefficients related to an elliptic contour. For this contour, the conformal mapping reduces to $\xi(w)=w+\gamma_{0}+\gamma_{1}w^{-1}$, where $\gamma_{0}$ is the centre of the ellipse and $\gamma_{1}=1-b$, with $b$ the major semiaxis. This maximises the memory efficiency of the algorithm, as the recurrence relation in Eq. 49 only depends on the two previous polynomials,
[…]
The coefficients $c_{n}$ of the Faber series presented in the main text (Eq. 10) are straightforwardly computed by performing the contour integral in Eq. 51 for the function $f(z)=e^{-i\delta t_{s}z}$. This is done through the use of the identity [74],
@@
[…]
@@ Appendix B General Features on the Hatano-Nelson Model
In this Appendix, we review the essential characteristics of the non-interacting Hatano-Nelson model, highlighting the importance of boundary conditions on the resulting physical properties. Specifically, we examine the diagonalisation of the non-interacting Hatano-Nelson model (Eq. 18) under both open and periodic boundary conditions. For OBCs, the Hamiltonian can be diagonalised through a similarity transformation. This is done using the $\text{GL}\left(1\right)$ gauge transformation[34],
@@
$\mathcal{H}_{\text{HN}}=-\sum_{n=0}^{L-2}\left[\frac{e^{\theta}\left(J+\gamma\right)}{2}p_{n}^{\dagger}q_{n+1}+\frac{e^{-\theta}\left(J-\gamma\right)}{2}p_{n+1}^{\dagger}q_{n}\right]$ (57)
@@
Unlike the previous case where the eigenstates were confined to the chain ends, under periodic boundary conditions both the right and left eigenstates are indistinguishable and delocalised. Furthermore, the spectra is always complex for all values of $\gamma$.
@@ Appendix C Comparison with MPS based Methods
where the Hamiltonian in Eq. 32 is decomposed as a sum of two body terms either acting on the left or right sites, $\mathcal{H}=\mathcal{H}_{\rm odd}+\mathcal{H}_{\rm even}$. Whilem in the first-order MPO, we use an Euler expansion of the time-evolution operator, $e^{-i\delta t\mathcal{H}}\simeq 1-i\delta t\mathcal{H}$, and represent the Hamiltonian through an MPO.
MPS calculations could be effective for this problem, as the entanglement entropy increases logarithmically over time and the domain-wall order is favoured by non-Hermiticity. However, the TEBD and MPO algorithms face errors that the Faber algorithm avoids. Firstly, there is a restriction on the maximum allowed time step. In the TEBD algorithm, the Hamiltonian is decomposed into smaller two-body terms that can be exactly exponentiated, as shown in Eq. 64. The approximation neglects the nonzero commutator term $\left[\mathcal{H}_{\rm odd},\mathcal{H}_{\rm even}\right]$, causing an error of order $\delta t^{2}$. To minimise this error, the time step must be small relative to the problem’s energy scale. This is evident in the left panel of Fig. 13, where longer integration times with a larger time step deviate from results obtained with Faber polynomials. A comparable error is present in the MPO algorithm due to the use of a first-order expansion. The Faber algorithm circumvents this issue by precisely representing the time evolution operator for any time step with any desired accuracy, though it requires the computation of the appropriate number of polynomials. Another limitation of the TEBD algorithm is the handling of long-range interactions with the Suzuki-Trotter decomposition, which requires the use of advanced techniques such as swap gates [124]. The TEBD and MPO techniques also face inaccuracies because of the truncation of the MPS bond dimension, a problem that the Faber algorithm avoids at the cost of working with a state encompassing the entire Hilbert space dimension. Of course, using the Faber polynomial limits the analysis to smaller system sizes. However, it can be advantageous in situations where MPS calculations fail, such as when the entanglement entropy scales proportionally to the system size.

### v3

@@ 7 Conclusions
This study offers additional insights into the entanglement transition in quantum spin chains exhibiting Non-Hermitian or Liouvillian skin effects. In Model A, we found that the area law behaviour of entanglement entropy remains for any nonzero $\gamma$, similar to the no-click limit. However, the sources are different: in the no-click limit, the area law stems from the single-particle skin effect, while in the monitored stochastic trajectory, it is due to the quantum jumps that relax the system into the fully polarised down spin-state, a product state. On the other hand, the dynamics with two-body jump operators (Model B), $L_{\ell}\propto S^{+}_{\ell}S^{-}_{\ell+1}$, allow for a non-equilibrium steady state with a magnetisation profile resembling the no-click limit of the Hatano-Nelson model. Furthermore, the average entanglement entropy still follows an area law for finite values of the ratio $\gamma/J$ in conditional dynamics. The measurement apparatuses effectively disentangle the state, suppressing the volume law that would otherwise be generated in unitary dynamics. This work extends previous studies that focused on the one-particle sector [108, 34].
Overall, our results support the utility and applicability of Faber polynomials in various research domains. This encompasses investigations into measurement-induced phase transitions, open quantum systems, and the exploration of purely non-unitary dynamics governed by non-Hermitian Hamiltonians. For instance, Faber Polynomials could be employed to study non-Hermitian Floquet problems in the high-frequency limit, where the Floquet Hamiltonian is well-defined and time-independent. Moreover, these polynomials can also be used to calculate general spectral properties of non-Hermitian Hamiltonians, potentially replicating the role of Chebyshev polynomials within the Kernel polynomial method [67]. This would complement the existing Non-Hermitian Kernel polynomial method [61, 62, 63]. Furthermore, the Faber polynomial approach could potentially be combined with MPS, similar to the developments already made in the Hermitian case [117, 118, 119, 120].
@@ Appendix A Further Details on Faber Polynomials
$F_{n+1}(z)=zF_{n}(z)-\sum_{j=0}^{n}\gamma_{j}F_{n-j}(z)-n\gamma_{n},\quad n>0,$ (49)
where $F_{0}(z)=1$. For our purposes, we are interested in using the Faber polynomials to approximate a given function of our non-Hermitian Hamiltonian, $f\left(\mathcal{H}\right)$. The domain $\mathcal{D}$, is defined by the spectrum of the Hamiltonian. Using Eq. (47), the expression in a Faber series is given by
$f(z):=\sum_{k=-\infty}^{+\infty}c_{k}F_{k}(z),$ (50)
@@
As stated in the main text, we perform this integral with $\rho=1$ by properly rescaling the Hamiltonian. Furthermore, we compute the Faber coefficients related to an elliptic contour. For this contour, the conformal mapping reduces to $\xi(w)=w+\gamma_{0}+\gamma_{1}w^{-1}$, where $\gamma_{0}$ is the centre of the ellipse and $\gamma_{1}=1-b$, with $b$ the major semiaxis. This maximises the memory efficiency of the algorithm, as the recurrence relation in Eq. (49) only depends on the two previous polynomials,
$F_{0}(z)$ $=1,$ (52)
$F_{1}(z)$ $=z-\gamma_{0},$
$F_{2}(z)$ $=\left(z-\gamma_{0}\right)F_{1}(z)-2\gamma_{1}F_{0}(z),$
$F_{n+1}(z)$ $=\left(z-\gamma_{0}\right)F_{n}(z)-\gamma_{1}F_{n-1}(z),\;n\geq 2.$
The coefficients $c_{n}$ of the Faber series presented in the main text (Eq. (10)) are straightforwardly computed by performing the contour integral in Eq. (51) for the function $f(z)=e^{-i\delta t_{s}z}$. This is done through the use of the identity [77],
@@
$F_{1}(z)$ $=T_{1}\left(\dfrac{z-\gamma_{0}}{2}\right),$ (55)
$F_{2}(z)$ $=T_{2}\left(\dfrac{z-\gamma_{0}}{2}\right)-1,$
$F_{n}(z)$ $=T_{n}\left(\dfrac{z-\gamma_{0}}{2}\right),n\geq 2.$
@@ Appendix B General Features on the Hatano-Nelson Model
In this Appendix, we review the essential characteristics of the non-interacting Hatano-Nelson model, highlighting the importance of boundary conditions on the resulting physical properties. Specifically, we examine the diagonalisation of the non-interacting Hatano-Nelson model (Eq. (18)) under both open and periodic boundary conditions. For OBCs, the Hamiltonian can be diagonalised through a similarity transformation. This is done using the $\text{GL}\left(1\right)$ gauge transformation[34],
@@
$\mathcal{H}_{\text{HN}}=-\sum_{n=0}^{L-2}\left[\frac{e^{\theta}\left(J+\gamma\right)}{2}p_{n}^{\dagger}q_{n+1}+\frac{e^{-\theta}\left(J-\gamma\right)}{2}p_{n+1}^{\dagger}q_{n}\right].$ (57)
@@
Unlike the previous case where the eigenstates were confined to the chain ends, under periodic boundary conditions, both the right and left eigenstates are indistinguishable and delocalised. Furthermore, the spectra is always complex for all values of $\gamma$.
@@ Appendix C Comparison with MPS based Methods
where the Hamiltonian in Eq. (32) is decomposed as a sum of two body terms either acting on the left or right sites, $\mathcal{H}=\mathcal{H}_{\rm odd}+\mathcal{H}_{\rm even}$. While in the first-order MPO, we use a Euler expansion of the time-evolution operator, $e^{-i\delta t\mathcal{H}}\simeq 1-i\delta t\mathcal{H}$, and represent the Hamiltonian through an MPO.
MPS calculations could be effective for this problem, as the entanglement entropy increases logarithmically over time and the domain-wall order is favoured by non-Hermiticity. However, the TEBD and MPO algorithms face errors that the Faber algorithm avoids. Firstly, there is a restriction on the maximum allowed time step. In the TEBD algorithm, the Hamiltonian is decomposed into smaller two-body terms that can be exactly exponentiated, as shown in Eq. (64). The approximation neglects the non-zero commutator term $\left[\mathcal{H}_{\rm odd},\mathcal{H}_{\rm even}\right]$, causing an error of order $\delta t^{2}$. To minimise this error, the time step must be small relative to the problem’s energy scale. This is evident in the left panel of Fig. 13, where longer integration times with a larger time step deviate from results obtained with Faber polynomials. A comparable error is present in the MPO algorithm due to the use of a first-order expansion. The Faber algorithm circumvents this issue by precisely representing the time evolution operator for any time step with any desired accuracy, though it requires the computation of the appropriate number of polynomials. Another limitation of the TEBD algorithm is the handling of long-range interactions with the Suzuki-Trotter decomposition, which requires the use of advanced techniques such as swap gates [126]. The TEBD and MPO techniques also face inaccuracies because of the truncation of the MPS bond dimension, a problem that the Faber algorithm avoids at the cost of working with a state encompassing the entire Hilbert space dimension. Of course, using the Faber polynomial limits the analysis to smaller system sizes. However, it can be advantageous in situations where MPS calculations fail, such as when the entanglement entropy scales proportionally to the system size.